OpenStudy (anonymous):

The function d(x) = 3x + 2 determines how many donuts need to be purchased for a meeting, where x is the number of people at the meeting. The office manager uses h(d(x)) to find the amount of money to bring for the donut purchase. The function h(x) = 2x + 7. Solve for how much money to bring when there are 8 people in the meeting. 25 26 52 59

OpenStudy (anonymous):

If there are 8 people in the meeting, you can calculate first how many donuts need to be purchased by using d(8)=3*8+2

OpenStudy (anonymous):

26?

OpenStudy (anonymous):

Yep. Then, to calculate how much money you'll need... h(d(x))= h(26) = 2*26+7

OpenStudy (anonymous):

59? :)

OpenStudy (anonymous):

I think so :)